# Agent-Based Insight into Eco-Choices: Simulating the Fast Fashion Shift

### v1

@@ V Model settings
Social media has two parameters to work with: the percentage of agents that get exposed to social media (determined by $\sigma\in[0.05,0.5]$) and social media bias ($\beta$). We study both, in particular in combination with increased communication. Models in this set are denoted by letter B.
@@
Every model is characterized by a set of parameters. The parameters and their definitions can be found in Table III. Please refer to the Appendix for a more detailed explanation of these parameters.
@@
TABLE III: Parameters used in the models
VI Results
In this section, we outline the settings of the models used in this project and present the results. The section is divided into three parts, each focusing on one mechanism of influence and its interaction with other influences.
VI-A Set A: peer pressure
[…]
We begin by looking at the communication levels and the effect they have on concerns. For the non polarized agentset, we run a baseline model with different levels of communication. We find that increased communication positively impacts the net changes in individuals’ average concerns, supporting previous research findings [25], [43], [21] Peer pressure is essential for behavioral change, as behaviors are not solely based on individual preferences but are reinforced by social expectations. Public willingness to accept, support, and actively participate in social, cultural, economic, and political changes is crucial for effective mitigation efforts. The graph showing changes in the probability of purchasing fast fashion can be found in Fig (1), while more information on changes in concerns can be found in Fig. (14) in the Appendix.
[…]
We conducted the same experiment with a polarized agent set. Similar to the non-polarized case, we observed an overall improvement in population concerns with increased communication levels. Interestingly, a medium level of communication ($\delta=0.3$) had a more positive effect on the non-polarized agent set than on the polarized one, whereas the highest communication level ($\delta=0.5$) had the opposite effect. The graph of changes in probabilities to purchase fast fashion is depicted in Fig. (1), while changes in concerns can be found in Fig. (15) in the Appendix.
Notably, in both simulations, the changes in environmental concerns are negative across all communication levels. This may be due to the initially high average environmental concerns at the start of the simulation. However, this high level of concern does not translate into corresponding actions. Since concerns are influenced by both opinions and behaviors, the general lack of action reduces overall concerns because peers’ behaviors impact normative expectations as much as, if not more than, their opinions. Moreover, the variability in opinions increases with the increase of the value of $\delta$, which expected given the polarizing nature of the agentset. (Fig. (16), Appendix.)
@@ VI-A Set A: peer pressure
[…]
Additionally, we studied the different levels of tolerance within the polarized agent set and their impact on communication. As the tolerance value ($\tau$) approaches 0.5, the agent set becomes less polarized, effectively becoming non-polarized when $\tau$ reaches its maximum value. We found that very high or very low tolerance levels are not constructive for communication (Fig. (17), Appendix.) Specifically, a tolerance level of $\tau=0.15$ yielded the best results in terms of desired changes. This suggests that extremely low tolerance ($\tau=0.05$) is detrimental to change, as it makes agents more reluctant to alter their opinions.
@@
VI-B Set B: social media influence
@@
[…]
@@
We begin by isolating the effects of social media on a non-polarized agent set, assuming no bias and focusing on exposure levels. We find that medium exposure ($\sigma=0.3$) has the most positive effect on the changes in concerns (Fig. (18), Appendix.) When exposure to social media is too low, not enough agents are influenced, leading to minimal change. Conversely, high exposure levels polarize agents due to the nature of the social media feedback loop function, increasing the number of anti-sustainability agents and reducing the average concern. This occurs because a significant portion of agents initially have high concerns, causing many to quickly polarize towards pro-sustainability. But higher exposure impacts those who may become polarized in the opposite direction, increasing the number of anti-sustainability agents and hence bringing the average concerns down.
[…]
@@
Next, we look at the different levels of biases on a non-polarized, ranging from very anti-sustainability ($\beta=-0.30$) to very pro-sustainability ($\beta=0.30$). We keep social media engagement constant (). We find that social media bias significantly impacts the effectiveness of communication in changing concerns. First, there is little difference in results between moderately pro-sustainability stances ($\sigma=0.15$ and $\sigma=0.30$), suggesting that even a moderate pro-sustainability bias can achieve a positive impact. However, moderately and strongly anti-sustainability stances ($\sigma=-0.15$ and $\sigma=-0.30$) hinder progress considerably (Fig. 19, Appendix). Social media often promotes increased consumption, especially in clothing, through in-platform shops, algorithms favoring shopping hauls, ads, and the ability to tag clothing in posts. This pro-consumption bias poses a significant threat to sustainable shopping habits, trapping individuals in a cycle of overconsumption. Conversely, if platforms favor sustainable content, they can promote more sustainable shopping habits [33], [41], [47].
[…]
@@
Moreover, the strong influence of social media persists in non-polarized agent sets with increased communication and social media exposure. In these cases of moderate influence, there are significant differences in overall changes, as shown in Fig. (3). The associated variances are illustrated in Fig. (20) in the Appendix, where variability in opinions is significantly larger with pro-sustainability social media bias compared to pro-consumption bias. This might be due to the initially high average probability of buying fast fashion at the start of the simulation. From these simulations, we learn that increased communication between agents is unlikely to combat social media, and significant influence of platforms remains dominant.
@@
[…]
For the polarized agent set, we examined the same scenario of different levels of exposure to social media, but with initially increased communication. We compared different levels of social media engagement. We found that social media has a smaller impact on overall changes in concerns compared to the non-polarized population (Fig. 21, Appendix). Additionally, lower levels of engagement ($\sigma=0.1$) positively impact concerns, which contrasts with the findings for the non-polarized agent set. On the contrary, medium ($\sigma=0.3$) higher levels ($\sigma=0.5$) of engagement had similar effects and impacted concerns and probability to buy fast fashion to a smaller extent. This was not the case for the non polarized agentset, highlighting how the same influence can impact agents differently depending on the nature of their communication.
(B5) $\delta=0.1$, $\tau=0.15$, $\sigma=0.35$ $\beta=-0.30,-0.15,0.00,0.15,0.30$ We observe a smaller impact of biases on the polarized agentset, compared to the non-polarized agent set, as shown in Fig. (4). This impact remains modest even with increased social media engagement. However, positive social media biases increase variability in all concerns and opinions, unlike negative biases, which show less impact on variability. By comparing Fig. (3) and Fig. (4), we see that social media biases affect the non-polarized agent set more significantly than the polarized agent set. This difference highlights how social media influences populations based on their preexisting views and the strength of their opinions.
@@ VI-C Set C: government interventions
[…]
We start by examining how state interventions influence shifts in average opinions within non-polarized agentsets. In our model, the state’s stance, represented by the value of $\zeta$, ranges from $\zeta=0.5$ (anti-sustainability) to $\zeta=1.5$ (pro-sustainability), reflecting different overall goals. Our initial focus is on the effects of government stances with increased communication levels. This choice is based on research suggesting that once a topic becomes politically or religiously prominent, it is likely to become polarizing and provoke increased communication [39]. Consequently, we limit our study to scenarios with elevated peer influence. We observe that agents are significantly influenced by government campaigns and stances, as expected in a non-polarized context. Interestingly, there is little difference in the outcomes between moderately pro-sustainability ($\zeta=1.2$) and highly pro-sustainability ($\zeta=1.5$) campaigns, as seen in Fig. (5). This suggests that a moderately pro-sustainability stance is sufficient for effective results, without needing an extreme position. Additionally, the probabilities of purchasing fast fashion cluster similarly for both moderately pro-fast fashion ($\zeta=0.8$) and pro-sustainability ($\zeta=1.2$) states, as shown by the individual distributions at the end of both simulations in the left histogram of Fig. (6). The clustering does occur around slightly different values, but the overall shape of the distribution is very similar.
@@
(C2) $\delta=0.4$, $\tau=0.15$, $\sigma=0.10$ $\beta=0$, $\zeta=0.5,0.8,1.0,1.2,1.5$
@@
(C3) $\delta=0.4$, $\tau=0.15$, $\sigma=0.40$ $\beta=-0.30,0.00,+0.30$, $\zeta=1.2$
The next question we explore is the role of social media in the policy-making process and its impact on progress. We previously noted that social media often exhibits pro-fast fashion biases, so we investigate how these and other stances influence shifts in shopping habits. We examine pro-fast fashion, neutral, and pro-sustainability social media biases within a polarized agent set, considering increased communication and social media engagement with a pro-sustainability state. Our simulations show that social media plays a significant role in shaping agents’ shopping habits (Fig. (23), Appendix.) It can either hinder or enhance progress, highlighting the importance for policy-makers to account for social media’s influence on the population.
(C4) $\delta=0.4$, $\tau=0.10,0.20,0.30$ $\sigma=0.10$, $\beta=0$, $\zeta=1.2$
Realistically, campaigns cannot continue indefinitely due to financial constraints and the diminishing impact of repeated exposure, a phenomenon known as “campaign fatigue” [40]. To address this, we investigate the effects of halting moderate pro-sustainability campaigns ($\zeta=1.2$) mid-simulation (around 250 time steps). Our focus is on a polarized agent set with varying levels of polarization. Given that higher tolerance reduces perceived controversy, we explore how different levels of polarization affect campaign effectiveness, using three tolerance levels with increased communication levels. We find that greater tolerance (lower polarization, $\tau=0.30$) results in more enduring effects, while the lowest tolerance ($\tau=0.10$) leads to a significant decay in concerns, often returning to pre-campaign levels (Fig. (24), Appendix.) In contrast, the non-polarized ($\tau=0.5$) agentset shows less promising results, where the achieved awareness levels fall to pre-campaign levels. Our findings suggest that a slightly polarized agent set is most likely to sustain the effects of a campaign over time.
(C5) $\delta=0.4$, $\tau=0.15$, $\sigma=0.10$ $\beta=-0.30,+0.30$, $\zeta=1.2$
@@ VII Conclusion
Our findings from models with increased socialization reinforce previous research findings on the importance of opinion sharing for fostering acceptance and supporting public conversations on challenging topics [25], [43].
@@
One goal of this study was to establish sensible measures for government implementation based on their objectives and the characteristics of the society in question. One of the key findings is that the government does not need to adopt an extremely proactive stance to achieve optimal results. We discovered that the state’s influence on public opinion reaches a point where further interventions yield insignificant returns. This indicates that more interventions are not necessarily better, since there is a limit to how much the state can affect its population.
@@ Acknowledgment
We would like to thank Dr. Silvia Blas Riesgo, for generously sharing her database on Spanish consumers. We would also like to thank the Spanish Department of Education for their financial support through the “Beca de Colaboración.”
@@ Appendix A Appendix
A-A Decision-making model.
@@
    Perceived Consumer Effectiveness: according to the Theory of Planned Behavior, Perceived Consumer Effectiveness influences purchasing power. However, empirical data revealed no correlation between perceived effectiveness regarding the impacts of shopping sustainably and individuals’ likelihood of purchasing it. Interestingly, nearly everyone expressed high levels of agreement with the following statements:
@@
    Even though most individuals agreed with these statements, it did not predict whether they shop sustainably or not.
@@
    Income: Interestingly, income played no role in people’s shopping habits. Furthermore, it did not predict shopping frequency or the average amount individuals spend on clothes. A linear regression analysis revealed that income only accounted for $1.9\%$ of the variation in monthly spending on clothing, and similar results were found for shopping frequency. This can be attributed to the nature of the fashion industry, particularly the fast fashion segment. The consistent decline in garment prices, despite inflation, has been a historical trend [18]) Additionally, people now allocate a smaller portion of their income to clothing than in the past [28]. While in the past, individual income significantly influences household expenditures on clothing [10], and some recent studies suggest this continues to be the case [17], it doesn’t seem to hold true for the Spanish consumer market, as there was no correlation between income and monthly expenditure on clothing.
@@
A-B Agent-Based Model: additional details.
A-B1 Polarized agentset - peer interaction function

### v2

@@ 5 Model settings
Social media has two parameters to work with: the portion of agents that get exposed to social media (determined by $\sigma\in[0.05,0.5]$), which is based on their levels of concerns) and social media bias ($\beta$). We study both, in particular in combination with increased communication. Models in this set are denoted by letter B.
@@
Every model is characterized by a set of parameters. The parameters and their definitions can be found in Table 3. Please refer to the Appendix for a more detailed explanation of these parameters.
@@
Table 3: Parameters used in the models
6 Results
In this section, we outline the settings of the models used in this project and present the results. These results are consistent across multiple runs, and although the reported numbers correspond to a single run, they are representative of the average outcomes. Additionally, the code and sample data — reflecting the original data’s distribution — are available on the CoMSES Net website. The section is divided into three parts, each focusing on one mechanism of influence and its interaction with other influences.
6.1 Set A: peer pressure
(A1) $\delta$: 0.1, 0.3, 0.5, $\tau=\text{N/A}$, $\sigma=0.1$, $\beta=0$, $\zeta=\text{N/A}$
We begin by looking at the communication levels and the effect they have on concerns. For the non polarized agentset, we run a baseline model with different levels of communication. We find that increased communication positively impacts the net changes in individuals’ average concerns, supporting previous research findings ([35], [56], [30]). Peer pressure is essential for behavioral change, as behaviors are not solely based on individual preferences but are reinforced by social expectations. Public willingness to accept, support, and actively participate in social, cultural, economic, and political changes is crucial for effective mitigation efforts. The graph showing changes in the probability of purchasing fast fashion can be found in Fig (1), while more information on changes in concerns can be found in Fig. (A.5) in the Appendix. Our findings for the non-polarized agentset are consistent with DeGroot’s model, where the average opinion converges to a consensus. However, it differs from the FJ model results, which postulates that an equilibrium is never reached. In our case, there are not enough stubborn agents and variability in initial opinions to reach results typical for an FJ model.
(A2) $\delta=0.1,0.3,0.5$, $\tau=0.15$, $\sigma=0.1$, $\beta=0$, $\zeta=\text{N/A}$
We conducted the same experiment with a polarized agent set. Similar to the non-polarized case, we observed an overall improvement in population concerns with increased communication levels. Interestingly, a medium level of communication ($\delta=0.3$) had a more positive effect on the non-polarized agent set than on the polarized one, whereas the highest communication level ($\delta=0.5$) had the opposite effect. The graph of changes in probabilities to purchase fast fashion is depicted in Fig. (1), while changes in concerns can be found in Fig. (A.6) in the Appendix.
Notably, in both simulations, the changes in environmental concerns are negative across all communication levels. This may be due to the initially high average environmental concerns at the start of the simulation. However, this high level of concern does not translate into corresponding actions. Since concerns are influenced by both opinions and behaviors, the general lack of action reduces overall concerns because peers’ behaviors impact normative expectations as much as, if not more than, their opinions. Moreover, the variability in opinions increases with the increase of the value of $\delta$, which expected given the polarizing nature of the agentset. (Fig. (A.7), Appendix.)
@@ 6.1 Set A: peer pressure
(A3) $\delta=0.1,0.3,0.5$, $\tau=0.15$, $\sigma=0.1$, $\beta=0$, $\zeta=\text{N/A}$
Additionally, we studied the different levels of tolerance within the polarized agent set and their impact on communication. As the tolerance value ($\tau$) approaches 0.5, the agent set becomes less polarized, effectively becoming non-polarized when $\tau$ reaches its maximum value. We found that very high or very low tolerance levels are not constructive for communication (Fig. (A.8), Appendix.) Specifically, a tolerance level of $\tau=0.15$ yielded the best results in terms of desired changes. This suggests that extremely low tolerance ($\tau=0.05$) is detrimental to change, as it makes agents more reluctant to alter their opinions. These results reinforce previous findings on how small changes to how individuals behave or how much they value similarity can shift collective dynamics.
@@
6.2 Set B: social media influence
@@
(B1) $\delta=0.1$, $\tau=\text{N/A}$, $\sigma=0.1,0.3,0.5$, $\beta=0$, $\zeta=\text{N/A}$
@@
We begin by isolating the effects of social media on a non-polarized agent set, assuming no bias and focusing on exposure levels. We find that medium exposure ($\sigma=0.3$) has the most positive effect on the changes in concerns (Fig. (A.9), Appendix.) When exposure to social media is too low, not enough agents are influenced, leading to minimal change. Conversely, high exposure levels polarize agents due to the nature of the social media feedback loop function, increasing the number of anti-sustainability agents and reducing the average concern. This is consistent with previous findings on mass media effects on cultural dynamics ([23]), which suggest that weak media influence promotes cultural homogeneity, while strong messaging sustains or enhances cultural diversity. This occurs because a significant portion of agents initially have high concerns, causing many to quickly polarize towards pro-sustainability. But higher exposure impacts those who may become polarized in the opposite direction, increasing the number of anti-sustainability agents and hence bringing the average concerns down.
(B2) $\delta=0.10$, $\tau=\text{N/A}$, $\sigma=0.10$, $\beta=-0.30,-0.15,0.00,0.15,0.30$. $\zeta=\text{N/A}$
@@
Next, we look at the different levels of biases on a non-polarized, ranging from very anti-sustainability ($\beta=-0.30$) to very pro-sustainability ($\beta=0.30$). We keep social media engagement constant (). We find that social media bias significantly impacts the effectiveness of communication in changing concerns. First, there is little difference in results between moderately pro-sustainability stances ($\sigma=0.15$ and $\sigma=0.30$), suggesting that even a moderate pro-sustainability bias can achieve a positive impact. However, moderately and strongly anti-sustainability stances ($\sigma=-0.15$ and $\sigma=-0.30$) hinder progress considerably (Fig. A.10, Appendix). Social media often promotes increased consumption, especially in clothing, through in-platform shops, algorithms favoring shopping hauls, ads, and the ability to tag clothing in posts. This pro-consumption bias poses a significant threat to sustainable shopping habits, trapping individuals in a cycle of overconsumption. Conversely, if platforms favor sustainable content, they can promote more sustainable shopping habits ([45], [55], [64]).
(B3) $\delta=0.40$, $\tau=\text{N/A}$, $\sigma=0.40$, $\beta=-0.15,+0.15$, $\zeta=\text{N/A}$
@@
Moreover, the strong influence of social media persists in non-polarized agent sets with increased communication and social media exposure. In these cases of moderate influence, there are significant differences in overall changes, as shown in Fig. (3). The associated variances are illustrated in Fig. (A.11) in the Appendix, where variability in opinions is significantly larger with pro-sustainability social media bias compared to pro-consumption bias. This might be due to the initially high average probability of buying fast fashion at the start of the simulation. From these simulations, we learn that increased communication between agents is unlikely to combat social media, and significant influence of platforms remains dominant.
@@
(B4) $\delta=0.4$, $\tau=0.15$, $\sigma=0.10,0.30,0.50$, $\beta=0$, $\zeta=\text{N/A}$
For the polarized agent set, we examined the same scenario of different levels of exposure to social media, but with initially increased communication. We compared different levels of social media engagement. We found that social media has a smaller impact on overall changes in concerns compared to the non-polarized population (Fig. A.12, Appendix). Additionally, lower levels of engagement ($\sigma=0.1$) positively impact concerns, which contrasts with the findings for the non-polarized agent set. On the contrary, medium ($\sigma=0.3$) higher levels ($\sigma=0.5$) of engagement had similar effects and impacted concerns and probability to buy fast fashion to a smaller extent. This was not the case for the non polarized agentset, highlighting how the same influence can impact agents differently depending on the nature of their communication.
(B5) $\delta=0.1$, $\tau=0.15$, $\sigma=0.35$, $\beta=-0.30,-0.15,0.00,0.15,0.30$
We observe a smaller impact of biases on the polarized agentset, compared to the non-polarized agent set, as shown in Fig. (4). This impact remains modest even with increased social media engagement. However, positive social media biases increase variability in all concerns and opinions, unlike negative biases, which show less impact on variability. By comparing Fig. (3) and Fig. (4), we see that social media biases affect the non-polarized agent set more significantly than the polarized agent set. This difference highlights how social media influences populations based on their preexisting views and the strength of their opinions.
@@ 6.3 Set C: government interventions
(C1) $\delta=0.4$, $\tau=\text{N/A}$, $\sigma=0.10$, $\beta=0$, $\zeta=0.5,0.8,1.0,1.2,1.5$
We start by examining how state interventions influence shifts in average opinions within non-polarized agentsets. In our model, the state’s stance, represented by the value of $\zeta$, ranges from $\zeta=0.5$ (anti-sustainability) to $\zeta=1.5$ (pro-sustainability), reflecting different overall goals. Our initial focus is on the effects of government stances with increased communication levels. This choice is based on research suggesting that once a topic becomes politically or religiously prominent, it is likely to become polarizing and provoke increased communication ([51]). Consequently, we limit our study to scenarios with elevated peer influence. We observe that agents are significantly influenced by government campaigns and stances, as expected in a non-polarized context. Interestingly, there is little difference in the outcomes between moderately pro-sustainability ($\zeta=1.2$) and highly pro-sustainability ($\zeta=1.5$) campaigns, as seen in Fig. (5). This suggests that a moderately pro-sustainability stance is sufficient for effective results, without needing an extreme position. Additionally, the probabilities of purchasing fast fashion cluster similarly for both moderately pro-fast fashion ($\zeta=0.8$) and pro-sustainability ($\zeta=1.2$) states, as shown by the individual distributions at the end of both simulations in the left histogram of Fig. (6). The clustering does occur around slightly different values, but the overall the shape of the distribution is very similar.
@@
(C2) $\delta=0.4$, $\tau=0.15$, $\sigma=0.10$, $\beta=0$, $\zeta=0.5,0.8,1.0,1.2,1.5$
@@
(C3) $\delta=0.4$, $\tau=0.15$, $\sigma=0.40$, $\beta=-0.30,0.00,+0.30$, $\zeta=1.2$
The next question we explore is the role of social media in the policy-making process and its impact on progress. We previously noted that social media often exhibits pro-fast fashion biases, so we investigate how these and other stances influence shifts in shopping habits. We examine pro-fast fashion, neutral, and pro-sustainability social media biases within a polarized agent set, considering increased communication and social media engagement with a pro-sustainability state. Our simulations show that social media plays a significant role in shaping agents’ shopping habits (Fig. (A.14), Appendix.) It can either hinder or enhance progress, highlighting the importance for policy-makers to account for social media’s influence on the population.
(C4) $\delta=0.4$, $\tau=0.10,0.20,0.30$, $\sigma=0.10$, $\beta=0$, $\zeta=1.2$
Realistically, campaigns cannot continue indefinitely due to financial constraints and the diminishing impact of repeated exposure, a phenomenon known as “campaign fatigue” ([52]). To address this, we investigate the effects of halting moderate pro-sustainability campaigns ($\zeta=1.2$) mid-simulation (around 250 time steps). Our focus is on a polarized agent set with varying levels of polarization. Given that higher tolerance reduces perceived controversy, we explore how different levels of polarization affect campaign effectiveness, using three tolerance levels with increased communication levels. We find that greater tolerance (lower polarization, $\tau=0.30$) results in more enduring effects, while the lowest tolerance ($\tau=0.10$) leads to a significant decay in concerns, often returning to pre-campaign levels (Fig. (A.15), Appendix.) In contrast, the non-polarized ($\tau=0.5$) agentset shows less promising results, where the achieved awareness levels fall to pre-campaign levels. Our findings suggest that a slightly polarized agent set is most likely to sustain the effects of a campaign over time.
(C5) $\delta=0.4$, $\tau=0.15$, $\sigma=0.10$, $\beta=-0.30,+0.30$, $\zeta=1.2$
@@ 7 Conclusion
Our findings from models with increased socialization reinforce previous research findings on the importance of opinion sharing for fostering acceptance and supporting public conversations on challenging topics ([35], [56]). In particular, large-scale shifts in public opinion depend considerably on the overall receptiveness of the broader population ([61]). This underscores the importance of cultivating environments where individuals are open to influence and engaged in peer-to-peer discussions, as collective change is more likely to emerge when many moderately connected people are socially primed to update their views.
@@
One goal of this study was to establish sensible measures for government implementation based on their objectives and the characteristics of the society in question. One of the key findings is that the government does not need to adopt an extremely proactive stance to achieve optimal results. We discovered that the state’s influence on public opinion reaches a point where further interventions yield insignificant returns. This indicates that more interventions are not necessarily better, since there is a limit to how much the state can affect its population. These results are consistent with existing studies that explore mechanisms to overcome public inaction and drive widespread pro-environmental behavior change ([29]).
@@
While our study provides valuable insights, it is based on several simplifying assumptions that could be refined in future research. For example, agents currently form friendships randomly and do not have the ability to end or create new ties based on opinion differences. Allowing such dynamics could lead to more realistic formations of homogeneous groups, reflecting real-world social behavior more closely. Similarly, interactions are modeled as unilateral and equally influential for all agents, whereas in reality, interactions tend to be bilateral with varying levels of influence among individuals. Additionally, agents interact on only one topic per time step, which complicates the direct correspondence between model time and real-world timelines. This relates to another limitation our study faces, which is the lack of longitudinal data necessary for the validation of the model.
[…]
Finally, the decision-making model is specifically calibrated to the Spanish market, which may limit the applicability of our results to other populations. Cultural differences and varying values placed on clothing both across and within countries mean that the model would need to be adjusted using local data to be relevant elsewhere.
@@ Acknowledgment
We would like to thank Dr. Silvia Blas Riesgo, for generously sharing her database on Spanish consumers.
This work was supported by the Ministry of Education, Vocational Training and Sports (MEFPD).
@@ Appendix A Appendix
A.1 Decision-making model.
@@
    Perceived Consumer Effectiveness: according to the Value-Belief-Norm (VBN) theory of environmental behavior ([54]), perceived consumer effectiveness influences purchasing power. However, our empirical data revealed no correlation between perceived consumer effectiveness regarding the impacts of shopping sustainably and individuals’ likelihood of purchasing it. Interestingly, nearly everyone expressed high levels of agreement with the following statements:
@@
    Even though most individuals agreed with these statements, it did not predict whether they shop sustainably or not. The study ([41]) also suggests that “sustainable awareness” is not strongly associated with sustainable purchasing behavior. Since definition of “sustainable awareness” combined our definitions of “Environmental concerns” and “Working conditions concerns,” it is not straightforward whether our findings agree, but we also find that environmental concerns have little impact on people’s purchasing behavior.
@@
    Income: Interestingly, income played no role in people’s shopping habits. Furthermore, it did not predict shopping frequency or the average amount individuals spend on clothes. A linear regression analysis revealed that income only accounted for $1.9\%$ of the variation in monthly spending on clothing, and similar results were found for shopping frequency. This can be attributed to the nature of the fashion industry, particularly the fast fashion segment. The consistent decline in garment prices, despite inflation, has been a historical trend ([25]). Additionally, people now allocate a smaller portion of their income to clothing than in the past ([38]). While in the past, individual income significantly influences household expenditures on clothing ([13]), and some recent studies suggest this continues to be the case ([24]), it doesn’t seem to hold true for the Spanish consumer market, as there was no correlation between income and monthly expenditure on clothing.
@@
A.2 Agent-Based Model: additional details.
A.2.1 Polarized agentset - peer interaction function